# Simultaneous Calibration of Noise Covariance and Kinematics for State Estimation of Legged Robots via Bi-level Optimization

### v2

@@ III-A2 Measurement Model
$\begin{bmatrix}G_{p}\\ G_{v}\end{bmatrix}\!=\!\begin{bmatrix}J(\alpha)&\mathbf{0}_{3\times 3}&\mathbf{0}_{3\times 3}\\ \dot{J}(\alpha,\dot{\alpha})+\omega^{\times}J(\alpha)&J(\alpha)&-[\,fk(\alpha)\,]_{\times}\end{bmatrix}\in\mathbb{R}^{6\times 9}.$
Thus, the induced measurement-noise covariances are $R_{\mathrm{pf}}\;=\;G_{p}\,R_{z}\,G_{p}^{\top}$, $R_{\mathrm{vf}}\;=\;G_{v}\,R_{z}\,G_{v}^{\top}$. If residuals terms $\mathbf{n}$ (8) are formed in $\mathrm{W}$ via $\mathbf{n}_{\mathrm{W}}=\mathbf{R}_{\mathrm{WB}}\;\mathbf{n}_{\mathrm{B}}$, the corresponding covariances becomes $R\mapsto\mathbf{R}_{\mathrm{WB}}\,R\,\mathbf{R}_{\mathrm{WB}}^{\top}$.
A binary contact flag $c\in\{0,1\}$, derived from pressure- or proprioceptive-based contact sensing, modulates the leg-odometry velocity channel: the base-velocity measurement is used only in stance ($c=1$) and suppressed in swing ($c=0$). In practice, the velocity channel’s contribution to the negative log-likelihood (objective) (III-B) (8) is multiplied by $c$.
III-B State Estimation as an MAP Problem
The optimal state estimator solves the Maximum A Posterior (MAP) problem:
$\hat{\mathbf{x}}^{*}$ $=\underset{\mathbf{x}_{[0,T]}}{\text{argmax}}\ \;{\mathbf{p}(\mathbf{x}_{[0,T]}|\mathbf{y}_{[0,T]})},$
$=\underset{\mathbf{x}_{[0,T]}}{\text{argmin}}\;\left(-\ln(\mathbf{p}(\mathbf{y}|\mathbf{x}))-\ln(\mathbf{p}(\mathbf{x}))\right),$ (7)
where the notation $(\cdot)_{[i,j]}$ denotes the sequence of vectors that starts from time index $i$ and ends at time index $j$. $\mathbf{p}(\mathbf{x}_{[0,T]}|\mathbf{y}_{[0,T]})$ is the posterior probability of the state trajectory $\mathbf{x}_{[0,T]}$, and $\hat{(\cdot)}^{*}$ denotes the optimal estimated solution.
For the estimation problem with the above-mentioned process and measurement models with additive Gaussian noises, the MAP problem can be solved via the Full information Estimation (FIE) formulation, with model descriptions formulated as state constraints:
$\min_{\mathcal{X}_{[0,T]}}\ \Gamma(\mathbf{x}_{0})$ $+\sum_{k=0}^{T-1}{\mathbf{w}_{k}^{\top}Q_{k}^{-1}\mathbf{w}_{k}}+\sum_{k=0}^{T}{\mathbf{n}_{k}^{\top}R_{k}^{-1}\mathbf{n}_{k}},$ (8)
$\text{s.t.}\quad\mathbf{x}_{k+1}$ $=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{w}_{k},\ \forall k=0,\ldots,T-1,$ (9)
$\mathbf{y}_{k}$ $=\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{n}_{k},\ \forall k=0,\ldots,T,$ (10)
$g(\mathbf{x}_{k})$ $\leq 0,\ \forall k=0,\ldots,T.$ (11)
where $\mathcal{X}_{[0,T]}$ is the state and noise trajectories up to $T$:
$\mathcal{X}_{[0,T]}$ $=\begin{bmatrix}\mathbf{x}_{[0,T]}^{\top}\;\mathbf{w}_{[0,T-1]}^{\top}\;\mathbf{n}_{[0,T]}^{\top}\end{bmatrix}^{\top}$,
$\mathbf{w}_{[0,T-1]}$ $=\begin{bmatrix}\mathbf{w}_{0}^{\top}&\mathbf{w}_{1}^{\top}&\cdots&\mathbf{w}_{T-1}^{\top}\end{bmatrix}^{\top}$,
$\mathbf{n}_{[0,T]}$ $=\begin{bmatrix}\mathbf{n}_{0}^{\top}&\mathbf{n}_{1}^{\top}&\cdots&\mathbf{n}_{T}^{\top}\end{bmatrix}^{\top}$.
$g(\cdot)$ denotes general state constraints. Given a Gaussian-distributed prior $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{x}_{\text{prior}},P_{0})$, the prior cost $\Gamma(x_{0})$ is defined as: $\Gamma(\mathbf{x}_{0})=(\mathbf{x}_{0}-\mathbf{x}_{\text{prior}})^{\top}P_{0}^{-1}(\mathbf{x}_{0}-\mathbf{x}_{\text{prior}})$.
@@ III-B State Estimation as an MAP Problem
The major drawback of FIE is that its computational cost grows with the horizon $T$, making online use practically intractable. MHE [kang24] instead uses a fixed past window of measurements and repeatedly solves the problem on a shifting horizon (as in MPC), yielding bounded and predictable cost. Because legged estimation is real-time, EKF variants with analytical updates are commonly used instead of windowed optimization. For offline calibration, we exploit all available data and therefore adopt FIE.
@@ IV Problem Statement
With properly selections of the noise covariance matrices $Q$ and $R$, the commonly applied estimators such as Extended Kalman Filters (EKFs) [bloesch2013state, rotella2014state], invariance EKFs [hartley2020contact, he2025invariant], and MHE [kang24] can yield satisfying estimates. However, it is widely known that extensive expert tuning of $Q$ and $R$ is inevitable on complex robots, and it can easily lead to sub-optimal, inconsistent or even diverging estimates on real robots. Additionally, inaccurate kinematics can further degrade estimate stability and precision. Therefore, we aim to provide an automatic and tuning-free method to optimally calibrate $Q$ and $R$ along with uncertain kinematics so that any reasonable choices of the estimators, regardless of EKFs or MHEs depending on the computational budget, can yield precise and consistent estimation on diverse set of tasks.
@@ IV-A Assumptions and Objective
Objectives: Our objective is thus to calibrate the optimal covariance matrices $Q$ and $R$ along with uncertain kinematic parameters so that the estimated results will be as close as possible to the ground truth using our process and measurement models. Without losing generality, for legged robots, we assume the last link on each leg has uncertain kinematic information; other linkage can be assumed to have uncertainties as well. This kinematic uncertainty on the last link (shin/calf) results in an uncertain foot position. Additionally, the motion capture device will capture markers that are placed on the torso. Therefore, besides uncertain shin kinematics, $\mathbf{x}_{\text{GT}}$ will not directly measure the robot state but with a constant kinematic offset, which will need to be calibrated along with the calibration of $Q$ and $R$.
@@ IV-B Bi-level Optimization
$\underset{\bm{\theta}\in\mathbb{R}^{m}}{\min}\quad$ $\mathcal{L}\bigl(\hat{\mathbf{x}}\left(\bm{\theta}\right),\mathbf{x}_{\mathrm{GT}},\bm{\theta}\bigr),$ (12)
@@
$\hat{\mathbf{x}}\left(\bm{\theta}\right)=\mathop{\mathrm{arg}\,\min}\limits_{\mathbf{x}\in\mathbb{R}^{n\times(T+1)}}\;\mathcal{J}\bigl(\mathbf{x},\bm{\theta}\bigr).$
@@
$p^{\mathrm{W}}_{\mathrm{real},j}\;=\;p^{\mathrm{W}}_{{\mathrm{foot}},j}\;+\;\mathbf{R}_{\mathrm{WC_{j}}}\,\bm{\theta}_{\mathrm{foot},j},$ (13)
in which the offset $\bm{\theta}_{\mathrm{foot},j}$ is rotated from $C_{j}$ to the world frame by $\mathbf{R}_{\mathrm{WC}_{j}}$ and added to the nominal foot position $p^{\mathrm{W}}_{\mathrm{foot},j}$ to yield the actual contact location $p^{\mathrm{W}}_{\mathrm{real},j}$. This time-invariant offset represents kinematic errors that come from mounting tolerances or manufacture inaccuracy. Note that, the observation vector $\mathbf{y}(\bm{\theta})$ (10) and the kinematic mapping matrices $\mathbf{G}(\bm{\theta})$ (III-A2) depend explicitly on $\bm{\theta}_{\mathrm{foot}}$.
At the lower level, we fix the vector $\bm{\theta}$ and solve the MAP problem, where $\mathcal{J}$ is the MAP cost in (III-B), (8). Given a covariance specification $\Sigma(\bm{\theta})$, where $\Sigma(\bm{\theta})$ is the collection of $Q$ and $R$, the solution of the MAP yields an estimated trajectory $\hat{\mathbf{x}}(\bm{\theta})\in\mathcal{X}$.
The estimated state at time $k$ in trajectory is defined as $\hat{\mathbf{x}}_{k}\;=\;(\mathbf{T}_{k},\,\mathbf{e}_{k})\;\in\;\mathrm{SE}(3)\times\mathbb{R}^{n_{\mathrm{eucl}}}$, where $\mathbf{T}_{k}$ denotes the base pose and $\mathbf{e}_{k}$ collects Euclidean sub-states such as base velocity and contact foot positions. The full trajectory lies on the manifold $\mathbf{X}\;\in\;\mathrm{SE}(3)^{T+1}\times\mathbb{R}^{n_{\mathrm{eucl}}(T+1)}$.
@@ IV-C Constraint Set $\mathcal{C}$
Given a positive-definite matrix $\Sigma_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$, we define its vectorization $\bm{\theta}_{\text{cov}}=\operatorname{vec}(\Sigma_{i})\in\mathbb{R}^{d_{i}(d_{i}+1)/2}$ by stacking the independent entries (upper triangular part including the diagonal) in a fixed order. During optimization, the parameter vector must remain in the constraint set $\mathcal{C}=\left\{\bm{\theta}\in\mathbb{R}^{m}\,\middle|\,\underline{\bm{\theta}}\leq\bm{\theta}\leq\overline{\bm{\theta}},\quad\Sigma_{i}(\bm{\theta})\in\mathbb{S}^{d_{i}}_{++},\;\forall i\right\}$, where $\underline{\bm{\theta}},\overline{\bm{\theta}}\in\mathbb{R}^{m}$ denote element-wise lower and upper bounds on each parameter, including both variances, correlations, and kinematic paramters. The set $\mathbb{S}^{d_{i}}_{++}$ denotes the cone of symmetric positive definite matrices of size $d_{i}\times d_{i}$.
@@ V Method
We now present our method to solve the bi-level optimization problem in (12). We adopt an iterative optimization approach: the lower-level problem (8) is solved using a standard interior-point method [wachter2006implementation], while the upper-level problem leverages its solution and Karush–Kuhn–Tucker (KKT) condition to update the parameters $\bm{\theta}$ via an adaptive Frank–Wolfe algorithm that iteratively refines the estimate until convergence. The index $t\in\mathbb{N}$ denotes the Frank–Wolfe iteration number.
@@
1:Input: $\mathcal{L}$, $\bm{\theta}^{t}$, $\mathcal{C}_{t}$, $t$
@@
subject to $\;\mathbf{s}\in\mathcal{C}_{t}\subset\mathbb{R}^{m}$
4:Step-size: $\gamma_{t}\!\in\!(0,1]\leftarrow\textsc{LineSearch}\!\left(\mathcal{L},\nabla\mathcal{L},\bm{\theta}^{t},\mathbf{s}^{\star}\right)$
5:Update: $\bm{\theta}^{t+1}\leftarrow(1-\gamma_{t})\,\bm{\theta}^{t}+\gamma_{t}\,\mathbf{s}^{t}$
@@ V-A Frank–Wolfe Algorithm for the Upper Level Optimization
Given that the constraint set $\mathcal{C}$ is convex and compact, we minimize the upper-level objective over $\mathcal{C}$ using the Frank-Wolfe algorithm, which preserves constraints without requiring projections. Given the upper-level gradient $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})$ computed as in Sec. (V-B2), each iteration solves a Linear Minimization Oracle (LMO) that minimizes the first-order approximation of $\mathcal{L}$ over $\mathcal{C}$, and updates the candidate solution by taking a convex combination of the current point with the LMO solution through line search.
To enhance numerical stability, each Frank–Wolfe subproblem is restricted to a local trust region around the current iterate $\bm{\theta}^{(t)}$: $\mathcal{C}_{t}=\mathcal{C}\cap\Bigl\{\bm{\theta}\in\mathbb{R}^{m}\;\big|\;\|\bm{\theta}-\bm{\theta}^{(t)}\|_{\infty}\leq\Delta_{t}\Bigr\}$, where $\Delta_{t}>0$ denotes the trust-region radius. This intersection preserves convexity and compactness, ensuring that the linear minimization oracle remains well-posed.
Linear Minimization Oracle: The LMO is implemented using two approaches:
@@
s.t. $\quad\underline{\mathbf{s}}\leq\mathbf{s}\leq\overline{\mathbf{s}},\;\;\sum_{i}s_{i}F_{i}\;\succeq\;\epsilon I.$
@@
Armijo Line Search: In practice, predefined decreasing step-size schedules often induce oscillations in $\mathcal{L}$ and make $\|\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})\|_{2}$ converge only after many iterations. We therefore adopt Algorithm 2 to choose an adaptive step size $\gamma_{t}$ that yields the sufficient decrease per iteration along the feasible FW direction returned by the LMO. We select $\gamma_{t}$ to satisfy
$\mathcal{L}\!\left(\bm{\theta}^{t}+\gamma_{t}\big(\mathbf{s}^{\star}-\bm{\theta}^{t}\big)\right)\leq\mathcal{L}(\bm{\theta}^{t})\;+\;\rho\,\gamma_{t}\;\big(\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})\big)^{\top}\big(\mathbf{s}^{\star}-\bm{\theta}^{t}\big),$
where $\rho\in(0,1)$ enforces a fixed fraction of the predicted decrease. The accepted $\gamma_{t}$ from Algorithm 2 is then used to update $\bm{\theta}$ in Algorithm 1.
Algorithm 2 Armijo Line Search (backtracking for $\gamma_{t}$)
1:Parameters: $\rho\in(0,1)$, $\beta\in(0,1)$, $\gamma\leftarrow 1$
2:while
$\mathcal{L}\!\left(\bm{\theta}^{t}+\gamma(\mathbf{s}^{\star}\!-\bm{\theta}^{t})\right)>\mathcal{L}(\bm{\theta}^{t})+\rho\,\gamma_{t}\big(\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})\big)^{\top}\!\big(\mathbf{s}^{\star}\!-\bm{\theta}^{t}\big)$
do $\gamma\leftarrow\beta\,\gamma$
3:end while
4:Return: $\gamma_{t}\leftarrow\gamma$

### v3

@@ III-A2 Measurement Model
$\mathbf{G}_{k}=\begin{bmatrix}\mathbf{J}(\tilde{\bm{\alpha}}_{k})&\mathbf{0}_{3\times 3}&\mathbf{0}_{3\times 3}\\ \mathrm{D}\mathbf{J}(\tilde{\bm{\alpha}}_{k})[\dot{\tilde{\bm{\alpha}}}_{k}]+[\tilde{\bm{\omega}}^{c}_{k}]_{\times}\mathbf{J}(\tilde{\bm{\alpha}}_{k})&\mathbf{J}(\tilde{\bm{\alpha}}_{k})&-[\mathrm{fk}(\tilde{\bm{\alpha}}_{k})]_{\times}\end{bmatrix}$. (3)
where $\tilde{\bm{\omega}}^{c}_{k}:=\tilde{\bm{\omega}}_{k}-\mathbf{b}_{\omega,k}$ and $\mathrm{D}\mathbf{J}(\bm{\alpha})[\mathbf{u}]$ denotes directional derivative of the matrix-valued map $\mathbf{J}$ at $\bm{\alpha}$ along $\mathbf{u}\in\mathbb{R}^{n_{\alpha}}$. This formulation explicitly capture coupling between IMU gyro noise and leg odometry uncertainty. Finally, a binary contact flag $c_{k}\in\{0,1\}$ gates the velocity residual, such that the effective information matrix is $\mathbf{\Omega}_{v}=c_{k}\mathbf{R}_{v,k}^{-1}$, ensuring velocity updates occur only during reliable stance phases.
III-B State Estimation as a Maximum A Posteriori Problem
We formulate the optimal state estimation problem within a Maximum A Posteriori (MAP) framework over a finite horizon $T$, the optimal state trajectory $\mathbf{x}_{0:T}\!:=\!\{\mathbf{x}_{0},\dots,\mathbf{x}_{T}\}$ is inferred by maximizing the posterior density given the sequence of inputs $\mathbf{u}_{0:T-1}$ and measurements $\mathbf{y}_{0:T}$:
$\hat{\mathbf{x}}_{0:T}^{\text{MAP}}$ $=\underset{\mathbf{x}_{0:T}}{\arg\max}\;p(\mathbf{x}_{0:T}\mid\mathbf{y}_{0:T},\mathbf{u}_{0:T-1})$
$=\underset{\mathbf{x}_{0:T}}{\arg\min}\left(-\ln p(\mathbf{y}_{0:T}\!\mid\!\mathbf{x}_{0:T})-\ln p(\mathbf{x}_{0:T}\!\mid\!\mathbf{u}_{0:T-1})\right).\vskip-4.0pt$
Under the Markov assumption and additive Gaussian noise models, it reduces to a deterministic Full Information Estimation (FIE). We formulate this as a constrained nonlinear optimization over the decision variables $\mathcal{Z}_{0:T}$:
$\min_{\mathcal{Z}_{0:T}}\quad$ $\big\|\mathbf{x}_{0}-\check{\mathbf{x}}_{0}\big\|_{\mathbf{P}_{0}^{-1}}^{2}+\sum_{k=0}^{T-1}\big\|\mathbf{w}_{k}\big\|_{\mathbf{Q}_{k}^{-1}}^{2}+\sum_{k=0}^{T}\big\|\mathbf{n}_{k}\big\|_{\mathbf{R}_{k}^{-1}}^{2},$
s.t. $\mathbf{x}_{k+1}=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{w}_{k},\quad\forall k\in\{0,\dots,T-1\},$
$\mathbf{y}_{k}=\mathbf{h}(\mathbf{x}_{k})+\mathbf{n}_{k},\quad\forall k\in\{0,\dots,T\},$
$\mathbf{g}(\mathbf{x}_{k})\leq\mathbf{0},\quad\forall k\in\{0,\dots,T\}.\vskip-5.0pt$ (4)
Here, the notation $\|\mathbf{v}\|_{\mathbf{\Sigma}^{-1}}^{2}\!:=\mathbf{v}^{\top}\mathbf{\Sigma}^{-1}\mathbf{v}$ denotes the squared Mahalanobis distance. The prior state distribution is given by $\mathbf{x}_{0}\sim\mathcal{N}(\check{\mathbf{x}}_{0},\mathbf{P}_{0})$. The decision variable vector $\mathcal{Z}_{0:T}$ explicitly includes the process noise $\mathbf{w}_{k}$ and measurement noise $\mathbf{n}_{k}$ as slack variables to satisfy the equality constraints:
$\mathcal{Z}_{0:T}:=\left[\mathbf{x}_{0:T}^{\top},\mathbf{w}_{0:T-1}^{\top},\mathbf{n}_{0:T}^{\top}\right]^{\top}.\vskip-3.0pt$
@@ III-B State Estimation as a Maximum A Posteriori Problem
$\mathbf{g}(\mathbf{x}_{k})\!\leq\!\mathbf{0}$ represents general state constraints (physical limits or obstacle avoidance). The major drawback of FIE is that its computational cost grows with the horizon $T$, making online use practically intractable. MHE [22] instead uses a fixed past window of measurements and repeatedly solves the problem on a shifting horizon (as in MPC), yielding bounded and predictable cost. Because legged estimation is real-time, EKF variants with analytical updates are commonly used instead of windowed optimization. For offline calibration, we exploit all available data and therefore adopt FIE.
@@ IV Problem Statement
With properly selections of the noise covariance matrices $\mathbf{Q}_{(\cdot)}$ and $\mathbf{R}_{(\cdot)}$, the commonly applied estimators such as Extended Kalman Filters (EKFs) [5, 30], invariance EKFs [15, 16], and MHE [22] can yield satisfying estimates. However, it is widely known that extensive expert tuning of $\mathbf{Q}_{(\cdot)}$ and $\mathbf{R}_{(\cdot)}$ is inevitable on complex robots, and it can easily lead to sub-optimal, inconsistent or even diverging estimates on real robots. Additionally, inaccurate kinematics can further degrade estimate stability and precision. Therefore, we aim to provide an automatic and tuning-free method to optimally calibrate $\mathbf{Q}_{(\cdot)}$ and $\mathbf{R}_{(\cdot)}$ along with uncertain kinematics so that any reasonable choices of the estimators, regardless of EKFs or MHEs depending on the computational budget, can yield precise and consistent estimation on diverse set of tasks.
@@ IV-A Assumptions and Objective
Objectives: Our objective is thus to calibrate the optimal covariance matrices $\mathbf{Q}_{(\cdot)}$ and $\mathbf{R}_{(\cdot)}$ along with uncertain kinematic parameters so that the estimated results will be as close as possible to the ground truth using our process and measurement models. Without losing generality, for legged robots, we assume the last link on each leg has uncertain kinematic information; other linkage can be assumed to have uncertainties as well. This kinematic uncertainty on the last link (shin/calf) results in an uncertain foot position. Additionally, the motion capture device will capture markers that are placed on the torso. Therefore, besides uncertain shin kinematics, $\mathbf{x}_{\text{GT}}$ will not directly measure the robot state but with a constant kinematic offset, which will need to be calibrated along with the calibration of $\mathbf{Q}_{(\cdot)}$ and $\mathbf{R}_{(\cdot)}$.
@@ IV-B Bi-level Optimization
$\underset{\bm{\theta}\in\mathbb{R}^{n_{\theta}}}{\min}\quad$ $\mathcal{L}\bigl(\hat{\mathbf{x}}\!\left(\bm{\theta}\right),\mathbf{x}_{\mathrm{GT}},\bm{\theta}\bigr),$ (5)
@@
$\hat{\mathbf{x}}\!\left(\bm{\theta}\right)=\!\mathop{\mathrm{arg}\,\min}\limits_{\mathbf{x}\in\mathbb{R}^{n_{x}\times(T+1)}}\;\mathcal{J}\bigl(\mathbf{x},\bm{\theta}\bigr).\vskip-6.0pt$
@@
$p^{\mathrm{W}}_{\mathrm{real},j}\;=\;p^{\mathrm{W}}_{{\mathrm{foot}},j}\;+\;\mathbf{C}_{\mathrm{WC_{j}}}\,\bm{\theta}_{\mathrm{foot},j},\vskip-2.0pt$ (6)
in which the offset $\bm{\theta}_{\mathrm{foot},j}$ is rotated from $C_{j}$ to the world frame by $\mathbf{C}_{\mathrm{WC}_{j}}$ and added to the nominal foot position $p^{\mathrm{W}}_{\mathrm{foot},j}$ to yield the actual contact location $p^{\mathrm{W}}_{\mathrm{real},j}$. This time-invariant offset represents kinematic errors that come from mounting tolerances or manufacture inaccuracy. Note that, the observation vector $\mathbf{y}(\bm{\theta})$ (2) and the kinematic mapping matrices $\mathbf{G}(\bm{\theta})$ (3) depend explicitly on $\bm{\theta}_{\mathrm{foot}}$.
At the lower level, we fix the vector $\bm{\theta}$ and solve the MAP problem, where $\mathcal{J}$ is the MAP objective in (4). Given a covariance specification $\Sigma(\bm{\theta})$, where $\Sigma(\bm{\theta})$ is the collection of $\mathbf{Q}_{(\cdot)}$ and $\mathbf{R}_{(\cdot)}$, the solution of the MAP yields an estimated trajectory $\hat{\mathbf{x}}(\bm{\theta})\in\mathcal{X}$. Following (1), we represent the estimated state as $\hat{\mathbf{x}}_{k}:=(\mathbf{T}_{k},\mathbf{e}_{k})\in\mathcal{G},\,\,\mathbf{X}:=(\hat{\mathbf{x}}_{0},\ldots,\hat{\mathbf{x}}_{T})\in\mathcal{G}^{T+1},$ where $\mathbf{T}_{k}\in\mathrm{SE}(3)$ is the base pose and $\mathbf{e}_{k}\in\mathbb{R}^{n_{\text{aug}}}$ stacks the Euclidean components in (1).
@@ IV-C Constraint Set $\mathcal{C}$
Given a positive-definite matrix $\Sigma_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$, we define its vectorization $\bm{\theta}_{\text{cov}}=\operatorname{vec}(\Sigma_{i})\in\mathbb{R}^{d_{i}(d_{i}+1)/2}$ by stacking the independent entries (upper triangular part including the diagonal) in a fixed order. During optimization, the parameter vector must remain in the constraint set $\mathcal{C}=\left\{\bm{\theta}\in\mathbb{R}^{n_{\theta}}\,\middle|\,\underline{\bm{\theta}}\leq\bm{\theta}\leq\overline{\bm{\theta}},\quad\Sigma_{i}(\bm{\theta_{\text{cov}}})\in\mathbb{S}^{d_{i}}_{++},\;\forall i\right\}$, where $\underline{\bm{\theta}},\overline{\bm{\theta}}\in\mathbb{R}^{n_{\theta}}$ denote element-wise lower and upper bounds on each parameter, including both variances, correlations, and kinematic paramters. The set $\mathbb{S}^{d_{i}}_{++}$ denotes the cone of symmetric positive definite matrices of size $d_{i}\times d_{i}$.
@@ V Method
We now present our method to solve the bi-level optimization problem in (5). We adopt an iterative optimization approach: the lower-level problem (4) is solved using a standard interior-point method [34], while the upper-level problem leverages its solution and Karush–Kuhn–Tucker (KKT) condition to update the parameters $\bm{\theta}$ via an adaptive Frank–Wolfe algorithm [19] that iteratively refines the estimate until convergence. The index $t\in\mathbb{N}$ denotes the Frank–Wolfe iteration counter, and superscript $(\cdot)^{t}$ denotes the value of variable at iteration $t$.
@@
1:Input: $\mathcal{L}$, $\bm{\theta}^{t}$, $\mathcal{C}^{t}$, $t$
@@
subject to $\;\mathbf{s}\in\mathcal{C}^{t}\subset\mathcal{C}$
4:Step-size: $\gamma^{t}\!\in\!(0,1]\leftarrow\textsc{LineSearch}\!\left(\mathcal{L},\nabla\mathcal{L},\bm{\theta}^{t},\mathbf{s}^{\star}\right)$
5:Update: $\bm{\theta}^{t+1}\leftarrow(1-\gamma^{t})\,\bm{\theta}^{t}+\gamma^{t}\,\mathbf{s}^{t}$
@@ V-A Frank–Wolfe Algorithm for the Upper Level Optimization
Given that the feasible set $\mathcal{C}$ (IV-C) is convex and compact, we minimize the upper-level objective over $\mathcal{C}$ using the Frank–Wolfe algorithm, which proceeds by optimizing a first-order (linear) surrogate of $\mathcal{L}$ at the current iterate.
With the upper-level gradient $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})$ computed as in Sec. (V-B2), each iteration solves the Linear Minimization Oracle (LMO) and updates the candidate solution via a convex combination of the current point and the LMO solution, with the step size chosen by line search. Equivalently, upper level optimization can be viewed as a sequential convex programming scheme: at iteration $t$, we minimize the linearization model of $\mathcal{L}$ over the convex set $\mathcal{C}^{t}$, yielding a convex LMO subproblem.
To enhance numerical stability, each Frank–Wolfe subproblem is restricted to a local trust region around the current iterate $\bm{\theta}^{t}$: $\mathcal{C}^{t}=\mathcal{C}\cap\Bigl\{\bm{\theta}\in\mathbb{R}^{n_{\theta}}\;\big|\;\|\bm{\theta}-\bm{\theta}^{t}\|_{\infty}\leq\Delta\!^{t}\Bigr\}$, where $\Delta\!^{t}>0$ denotes the trust-region radius. This intersection preserves convexity and compactness, ensuring that the linear minimization oracle remains well-posed.
Linear Minimization Oracle: The convex LMO can be solved using two alternative approaches:
@@
s.t. $\quad\underline{\mathbf{s}}\leq\mathbf{s}\leq\overline{\mathbf{s}},\;\;\sum_{i}s_{i}F_{i}\;\succeq\;\epsilon I.\vskip-10.0pt$
@@
Armijo Backtracking Step Size: Fixed, pre-defined decreasing step-size schedules can cause oscillations in $\mathcal{L}$ and slow the decay of $\|\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})\|_{2}$. We therefore use an Armijo backtracking line search to choose an adaptive step size $\gamma^{t}\in(0,1]$ along the feasible Frank–Wolfe direction $\mathbf{d}^{t}\triangleq\mathbf{s}^{\star}-\bm{\theta}^{t},$ returned by the LMO. Specifically, for given parameters $\rho\in(0,1)$ and $\beta\in(0,1)$, we initialize $\gamma\leftarrow 1$ and repeatedly shrink $\gamma\leftarrow\beta\gamma$ until the following sufficient-decrease condition is satisfied at $\gamma^{t}$
$\mathcal{L}\!\left(\bm{\theta}^{t}+\gamma^{t}\,\mathbf{d}^{t}\right)\leq\mathcal{L}(\bm{\theta}^{t})+\rho\,\gamma^{t}\,\big(\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})\big)^{\top}\mathbf{d}^{t}.\vskip-5.0pt$
The iterate is then updated as $\bm{\theta}^{t+1}=\bm{\theta}^{t}+\gamma^{t}(\mathbf{s}^{\star}-\bm{\theta}^{t}).$
For the convergence behavior, empirically with a good initialization, $\mathcal{L}$ behaves locally along the iterates as a convex $C^{1}$ function, so Frank–Wolfe with line search achieves $\mathcal{L}(\theta^{t})\!-\!\mathcal{L}^{\star}\!=\!O(1/t)$ [19]. Under poorer initialization, $\mathcal{L}$ is effectively nonconvex, instead the standard guarantee is $\min_{k\leq t}g(\theta^{k})\!=\!O(1/\sqrt{t})$. For the runtime, let $T_{\mathrm{FIE}}(T)$ denote the time of one lower-level FIE solve. At outer iteration $t$, line search tries $N_{\mathrm{bt}}^{t}$ step sizes and thus evaluates $\mathcal{L}(\cdot)$ $N_{\mathrm{bt}}^{t}$ times. When $T$ is large, the implicit-differentiation cost (V-B2) and the LMO cost are negligible compared to the lower-level solve, hence $T_{\mathrm{iter}}^{t}\!\!\!=\!(1+N_{\mathrm{bt}}^{t})\,T_{\mathrm{FIE}}(T)+O(T).$